40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 270 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  3759   Fri Oct 22 01:23:13 2010 KojiUpdatePSLFound the beat at 1064nm

[Koji / Suresh]

We worked on the 1064 beating a bit more.

- First of all, FSS and FSS SLOW servo were disabled not to have variating Xtal temp for the PSL.

- The PSL Xtal temp (T_PSL) was scanned from 22deg-45deg while we search the Xtal temp (T_Xend) for the Xend laser to have the beat freq well low (f<30MHz).
The pumping current for each laser was I_PSL = 2.101 [A] and I_Xend = 2.000 [A]

For a certain T_PSL, we found multiple T_Xend because the freq of the laser is not a monotonic function of the Xtal temperature. (see the innolight manual).

T_Xend to give us the beating was categorized in the three sets as shown in the figure. The set on "curve2" is the steadiest one. (Use this!)
The trends were quite linear but the slope was slightly off from the unity.

- T_PSL was scanned to see the trend of the PMC output.

The PMC was sometimes locked to the mode with lower transmission (V_PMCT ~ 3.0V).
When T_PSL ~ 31deg we consistently locked the PMC at higer transmission (V_PMCT ~ 5.3V).

At the moment we decided the operating point of T_PSL = 32.25 deg, V_PMCT = 5.34, where we found the beat at T_Xend=38.28deg.

- We cleaned up the PSL table more than how it was. Returned the tools to their original places.
The X-end laser was shut down and was left on the PSL table.

NEXT:
Kiwamu can move the X-end laser to the Xend and realign it.
Then we should be able to see the green beating quite easily.

Attachment 1: 101021_beat.pdf
101021_beat.pdf
  3758   Fri Oct 22 00:59:01 2010 yutaUpdateCDSnew MEDM screens, new filter banks(whitening and dewhitening related)

(Joe, Yuta)

Summary:
 No more discrimination for SIDE!
 We had all SIDE filters in SDSEN, so we splitted into SDSEN, SUSSIDE, and SDCOIL just like other DOFs. (Not SIDECOIL. SDCOIL from now on!)
 Also, there was a confusion in output filters, so we fixed the filter bank(dewhitening and 28HzELP).
 More than that, I found that "13HzCheby" in SUSPOS, SUSPIT, SUSYAW was actually doing ~1.6Hz Chebyshev, so I fixed it. (No wonder we had no damping when turing "13HzCheby" on!)
 Our new MEDM screen is Attachment #1.

Input filter design:
 Every optics, every OSEM has same analog whitening filter.
 So;

digital analog
30,100:3 30,100:3
ON OFF
OFF ON

Output filter design:
 For every SIDE coils and MC1, MC3 coils, they have analog 28Hz elliptic LPFs.
 So;
digital analog
28HzELP 28HzELP
ON OFF
OFF ON

 For MC2(and other main optics) UL/UR/LR/LL coils, they have analog dewhitening filters.
 So;
digital analog
InvDW SimDW DW
ON ON OFF
ON OFF ON

 InvDW and SimDW were in FM10 and FM9, but I swapped them to make it more consistent.
 Now, FM10 switches analog/digital output filiters for every optics.

  Let's put 28HzELP in FM9 so that FM9 switches analog/digital for every optics.

Quote:

Schematic:
 - whitening
   MC1 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC2 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC3 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
      D000210 has switches for bypassing analog WT(3,100:3). HIGH to bypass.
 - dewhitening
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,3 UL/UR/LR/LL coils
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,2,3 SIDE coils
  (SimDW)(InvDW) ... -> LSC Anti-imaging Board(D000186) -> Universal Dewhitening Board(D000183) -> MC2 UL/UR/LR/LL coils
     D000316 has switches for bypassing 28Hz elliptic LPF. HIGH to bypass.
     D000186 is 7570Hz elliptic LPFs.
     D000183 has switches for bypassing dewhitening filter. HIGH to bypass.
 See this wiki page for more comprehensive setup.

Next work:
 - Fix Simulink connection and logic so that analog/digital switching go correctly.
 - Check if the switching are correct for all channels by measuring transfer functions.
 - Currently, we are focusing on MCs, but we have to do the same fixing for other optics, too.
 - More than setting the right filter TF, there are switching settings like Zero History, Ramp, Zero Crossing...... We should investigate that.
 - Actually, TF of analog DW in D000183 doesn't look like the same with digital SimDW. Why??

Attachment 1: HappySideMEDM.png
HappySideMEDM.png
  3757   Thu Oct 21 21:35:16 2010 yutaUpdateCDSfixed whitening filter switches

(Joe, Yuta)

Summary:
 We found some mistakes in whitening filter switches yesterday(see elog #3748).
 One was the mis-connection between LL/LR and the second was the bit invert.
 So, we fixed it and checked that they are now switching correct by probing MAX333A.

Why we made mistake:
1. LL/LR mis-connection
  Simply, Simulink model was wrong.
  It occurred because UL/UR/LR/LL order is currently arbitrary and confusing.
  For now, we just swap the connection(Attached #1).
  In the future, we should fix all orders corresponding to the actual wiring order(UL > LL > UR > LR).

2. Bit invert
  It occurred from counter-intuitive output of Contec 32 BO(See Attachment #2(taken from this wiki page)).
  If "1," current goes from A to B, so MAX333A input will be 0V, which means analog WF is on.
  If "0," no current, so MAX333A input will be +15V, which means analog WF is off(bypassed).

Next work:
 - There are no digital WF in SDSEN inputs. So, we have to put them.
 - Dewhitening filters are totally messed up now. Switches are wrong, some optics have digital DW but some doesn't, some have 28Hz elliptic LPF.......
     We have to grand unify them.

Attachment 1: LLLR.png
LLLR.png
Attachment 2: Contec23BO.png
Contec23BO.png
  3756   Thu Oct 21 19:10:39 2010 AidanUpdatePSLFound the beat at 1064nm

Quote:

[Koji Suresh]

We found the beat at 1064nm. T(PSL)=26.59deg, T(X-end)=31.15deg.

The X-end laser was moved to the PSL table.

The beating setup was quickly constructed with mode matching based on beam profile measurements by the IR cards.
We used the 1GHz BW PD, Newfocus #1611-FS-AC.

As soon as we swept the Xtal temp of the X-end laser, we found the strong beating.

 

  3755   Thu Oct 21 18:45:50 2010 KojiUpdatePSLFound the beat at 1064nm

[Koji Suresh]

We found the beat at 1064nm. T(PSL)=26.59deg, T(X-end)=31.15deg.

The X-end laser was moved to the PSL table.

The beating setup was quickly constructed with mode matching based on beam profile measurements by the IR cards.
We used the 1GHz BW PD, Newfocus #1611-FS-AC.

As soon as we swept the Xtal temp of the X-end laser, we found the strong beating.

  3754   Thu Oct 21 15:59:28 2010 kiwamuUpdateLockingno green beat note : details

Last night I tried searching for a beat note signal with two different PD trans impedance gains.

Although I didn't find a  beat note signal.

 


 

- (1. trans impedance gain = 2400)

 I started with a trans impedance resistance of R=2.4k, which is 10 times bigger resistance than the original.

The total PD gain should be about 960 [V/W] theoretically if we assume the responsibility of the PD is 0.4 [A/W].

Then I checked the bandwidth of the RFPD using Jenne laser.

The bandwidth was about 30MHz, which is 3 times narrower than the original. And it agrees with our expectation.

As Koji and I mentioned at the last weekly meeting, the cut off frequency of an RFPD follows inverse square root of the trans impedance resistance R.

 

 bw.png

where C is a capacitance of the photo diode. (See this)

I was expecting the signal level of  -50 dBm / rtHz with a 23dB RF amplifier, assuming the line width of the signal is 10kHz.

 

 


- (2. trans impedance gain = 240) 

 I also tried it with the original trans impedance gain (see this entry). 

 R = 240 [Ohm]

 G = 96 [V/W]

 BW = 100 [MHz]  (I didn't measure it in this time)

 expected signal level = -70 dBm/rtHz


Quote:

 Still I didn't see any beat note signals..

 

  3753   Thu Oct 21 13:05:19 2010 KojiSummaryComputerselog change and rossa tex

This is cool though the projector is flashing the blue screen alternately.

I gave the dual head video card (ATI RADEON something) to Yuta a month ago.
It is on the top of Zita. This would make the things more fun.

Quote:

Yuta has also successfully set up zita to run the projector, so we should clear our some of the boxes and bookshelves in that area so that the projection can be larger.

 

  3752   Thu Oct 21 12:15:02 2010 ranaUpdatePEM6.9 Mag EQ in Gulf of California
Magnitude 6.9
Date-Time
Location 24.843°N, 109.171°W
Depth 10 km (6.2 miles) set by location program
Region GULF OF CALIFORNIA
Distances 105 km (65 miles) S of Los Mochis, Sinaloa, Mexico
125 km (75 miles) SW of Guamuchil, Sinaloa, Mexico
140 km (85 miles) NE of La Paz, Baja California Sur, Mexico
1200 km (740 miles) WNW of MEXICO CITY, D.F., Mexico
Location Uncertainty horizontal +/- 6.1 km (3.8 miles); depth fixed by location program
Parameters NST=187, Nph=187, Dmin=843.1 km, Rmss=1.17 sec, Gp=133°,
M-type=teleseismic moment magnitude (Mw), Version=6
Source
  • USGS NEIC (WDCS-D)
Event ID us2010crbl
  3751   Thu Oct 21 10:44:56 2010 steveMetaphysicsTreasure1987 supernova tapes plus....?

I'm cleaning out to make room for our new optical cabinet. Are we keeping these? There are  ~20  pieces of 10" od 1" wide tapes and large number of cassettes.

AJW,  Zucker,  Stuart A and Koji were notified in this matter.

Alan suggested to save data of Bruce Allen paper of observation of binary neutron stars in the 40m on 1994 November 14-20 and save back up tapes of his period in the 40m.

Mike: reels are not readable any more, it is time to let go

Attachment 1: P1060932.JPG
P1060932.JPG
  3750   Thu Oct 21 05:51:10 2010 kiwamuUpdateLockingno green beat note

 Still I didn't see any beat note signals..

 With a help from Suresh, Yuta and Rana, I tried searching for the green beat note by changing the temperature of the X end NPRO.

The noise level after it goes through an RF amplifier (G=23dB) was about -70dBm at 50MHz.

This noise may cover the beat note signal.

I am going to post some details later.

  3749   Thu Oct 21 01:11:48 2010 ranaSummaryComputerselog change and rossa tex

I deleted a bunch of categories from the elogd.cfg file. Not every kind of locking and activity gets its own activity. All of the things related to length sensing and control should be LSC. If there's a high pollen count its under PEM. etc., etc., etc.

I decided that the 'tetex' distribution which comes with CentOS is total BS and I removed it from rossa using 'yum erase tetex'. I installed TexLive using the instructions on the web; its much better, actually works, and also has RevTex 4.1 which allows Jenne and I to write our paper.

Eventually, we'll standardize our workstation setups.

Yuta has also successfully set up zita to run the projector, so we should clear our some of the boxes and bookshelves in that area so that the projection can be larger.

  3748   Wed Oct 20 21:43:25 2010 yutaUpdateCDSchecking whitening filters for MCs and messed up

(Joe, Yuta)

Background:
 We found that the damping servo for MC suspensions somehow worked when we turned off the 13Hz Chebyshev filters.
 But that does not meet our satisfaction, so we started checking every components.
 First of all is the whitening filters.
 If we turn on a digital whitening filter(WF), corresponding analog WF should turn off, and vice versa.

Reference:
 See DCC #D000210 for the analog circuit of WF. WF has 3Hz zero, 30Hz pole, 100Hz pole. MAX333A bypasses analog WF when supplied +15V(HIGH).

What we did:
 1. Compared the transfer function between MC2_SUSPOS_EXC and MC2_ULSEN_IN1 when digital WF on and off. When digital is on/off, analog should be off/on, so there should be difference but couldn't see.

 2. We went through the simulink model and found 2 mistakes in the logic. One is the conflict with other optics. Even if we turn on/off digital WF of MC2, it didn't switched analog WF of MC2. Two is the additional bit invert (but it turned out to be our misunderstanding).

 3. We (thought we) fixed it, rebuild it, and measured the TF again.(Attachment #1)

[Attached #1]
 The red/blue line is when digital WF is on/off. Blue should be bigger(+10dB @ 10Hz according to (analog) WTF) than red, but it was the opposite.

 4. To confirm that they are doing the opposite, I checked MAX333A input(pin#1,10,11,20) in "SUS PD Whitening Board" at 1X5-1-8B (which is for MC3) and found that switching is opposite. When I turned off/on the digital WF, MAX333A input was +15V/0V. It should be 0V/+15V.

 5. Also, I found that LLSEN digital WF switch switches LRSEN analog WF and vice versa.

[Attached #2]
 Transfer functions between MC2_SUSPOS_EXC and MC2_LRSEN_IN1 with;
   Red: LR digital off, LL digital on
   Blue: LR digital off, LL digital off
   Green: LR digital on, LL digital off
 As you can see, LL switch is the one which switches LR analog WF now.

Conclusion:
 Currently, digital WF on/off corresponds to analog WF on/off.
 Also, LL/LR digital WF switch is LR/LL analog WF switch now.


Next work:
 - fix the simulink models (or wiring)
 - check dewhitening filters

Schematic:
 - whitening
   MC1 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC2 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC3 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
      D000210 has switches for bypassing analog WT(3,100:3). HIGH to bypass.
 - dewhitening
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,3 UL/UR/LR/LL coils
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,2,3 SIDE coils
  (SimDW)(InvDW) ... -> LSC Anti-imaging Board(D000186) -> Universal Dewhitening Board(D000183) -> MC2 UL/UR/LR/LL coils
     D000316 has switches for bypassing 28Hz elliptic LPF. HIGH to bypass.
     D000186 is 7570Hz elliptic LPFs.
     D000183 has switches for bypassing dewhitening filter. HIGH to bypass.
 See this wiki page for more comprehensive setup.

Attachment 1: MC2ULSEN.png
MC2ULSEN.png
Attachment 2: MC2LRSEN.png
MC2LRSEN.png
  3747   Wed Oct 20 21:33:11 2010 KevinUpdatePSLQuarter Wave Plate Optimization

[Suresh and Kevin]

We placed the quarter wave plate in front of the 2W laser and moved the half wave plate forward. To make both wave plates fit, we had to rotate one of the clamps for the laser. We optimized the angles of both wave plates so that the power in the reflection from the PBS was minimized and the transmitted power through the faraday isolator was maximized. This was done with 2.1 A injection current and 38°C crystal temperature.

Next, I will make plots of the reflected power as a function of half wave plate angle for a few different quarter wave plate rotations.

  3746   Wed Oct 20 18:17:35 2010 Suresh, JenneUpdateSUSPRM assembly

We have positioned the guide rod and the wire-stand-off on the optic in the axial direction. 

We have selected six magnets whose magnetic strength is +/-5% of their mean strength (180 Gauss).  The measurement was made as follows:

1) each magnet was placed on its  end, on the top of a beaker held upside down. 

2) The Hall probe was placed directly under the magnet touching the glass from the other side (the inside of the beaker). 

This ensures that the relative position of the magnet and the probe remains fixed during a measurement.  And ensures that their separation is the same for each of the magnets tested. 

With this procedure the variation in the measured B field is less than +/- 10% in the sample of magnets tested.

  3745   Wed Oct 20 16:41:41 2010 steveOmnistructureGeneral laser chiller leaving for MIT
Attachment 1: P1060929.JPG
P1060929.JPG
  3744   Wed Oct 20 01:22:18 2010 yutaUpdateCDSfixed filters for MCs, but no damping

(Rana, Yuta)

Summary:

 The damping servo for MCs has not been working since Monday.
 We already found that some filter coefficients were wrong.(see elog #3742)
 So, we fixed it (just for MCs now).
 But still doesn't work.

What we did:
Fixed the filters for MC suspensions;
 1. In /cvs/cds/rtcds/caltech/c1/chans/ directory, we replaced C1MCS.txt with ./filter_archive/c1mcs/C1MCS_101019_101927.txt.
  This is the archive of the filter bank for MC suspensions, when filter designs were correct(when we hadn't opened it with foton yet).

 2. Deleted all the filter coefficients.
   By using emacs magical C-<space> C-x r k.

 3. Loaded with foton to get correct coefficients.

 4. Reloaded coefficients for C1MCS.

We now have the correct filters, but the damping still doesn't work.

 5. To confirm that the filters are now correct and the model is working correctly, we measured the transfer function between ULSEN input and ULCOIL output and compared with the calculated TF from the filter bank file (Attachment #1) .

 6. To calculate the designed function, we used the following matlab scripts;
    /cvs/cds/caltech/users/rana/mat/utilities/FotonFilter.m        # takes the foton file and returns the TF
    /cvs/cds/caltech/users/rana/mat/utilities/readFilterFile.m    # reads the foton file
    /cvs/cds/caltech/users/yuta/scripts/sentocoil.m                   # calculates the total TF from SEN to COIL

Result:
 As you can see from the attachment, filters seem fine now. (The red line is the measured and the blue line is the calculated function in the plot)
 But the damping (for POS, PIT, YAW) still doesn't work. Why????

Next work:

 - see if coils are pushing correctly
 - push cable connectors further!
 - check input and output filters (whitening, dewhitening)
 - when fixed, use my fully updated QAdjuster.py for adjusting Q for multiple channels automatically!!

Attachment 1: Screenshot.png
Screenshot.png
  3743   Tue Oct 19 22:37:28 2010 KojiUpdatePSLPSL table cleaning up

I cleaned up the scattered tools, optics, and mounts of the PSL table. I gathered those stuffs at the two coners.

At the end of the work I scanned the table with an IR viewer. (This is mandatory)
I put some beam block plates to kill weak stray beams.

One thing I like to call the attention is:

I found that some beam blocks were missing at around the PBSs just after the laser source.
Those PBSs tend to reject quite a lot of beam power
--- no matter how the HWPs/QWPs are arranged.

--- even at the backward side.
(remember that we have a faraday there.)

Particularly, there was no beam block at the forward rejection side of the first PBS where we dump the high power beam.

Be careful. 

  3742   Tue Oct 19 21:10:27 2010 yutaUpdateCDSbad filter coefficients for every suspensions!

(Koji, Yuta)

Summary:
 The damping servos of the MC suspensions has not been working since yesterday. They had worked on Friday.
 We found that some of the filter coefficients somehow changed from the correct value during dealing with the sampling frequency shift (from 2048Hz to 16384Hz). (see elog #3659)

What we did:
We thought that the problem occurred from the CDS update on Monday. So, we restored them using svn.
 1. Went to /opt/rtcds/caltech/c1/core/advLigoRTS and ran svn update using the revision number 2125.
   svn update -r 2125
 2. Rebuilt all the front ends.
   ssh -X c1sus
   cd /opt/rtcds/caltech/c1/core/advLigoRTS
   make clean-SYSNAME
   make SYSNAME
   make install-SYSNAME 
   (where SYSNAME=c1x02,c1sus,c1mcs,c1rms)
 3. Restarted the front end models.
   ssh -X c1sus
   cd /opt/rtcds/caltech/c1/scripts
   sudo startc1SYS 
   (where SYS=sus,mcs,rms)
 4. Restarted mx_streams.
   sudo /etc/restart_streams
See this wiki page for the restart procedures.

At this point we judged that the realtime system and "dataviewer" worked fine (by poking some suspensions / by measuring PSDs/TFs of the signals).

However, just restoring the RT system didn't fix the damping servos.

After that, we measured the transfer functions of the servo filters using DTT(diaggui on rosalba).
 5. The filter at MC1_SUSPOS, "3:0.0" should have a cut-off frequency at 3Hz, but it was around 0.3Hz.
 6. We used foton in order to look at the filter bank files (in /cvs/cds/rtcds/caltech/c1/chans). Then we found that the filter design was somehow changed to
   zpk([0],[0.375],2.66666,"n")
  instead of the correct one
   zpk([0],[3],0.333333,"n")

Why the filter designs changed?:
 We suspect that the filter designs were changed because of the replacement of sampling frequency in filter bank files(see elog #3659). We only replaced the lines like
  # SAMPLING SUS_MC3_SUSPOS 2048
 to
  # SAMPLING SUS_MC3_SUSPOS 16384
 and didn't changed the coefficients.
 This may caused a problem when opening the files with foton, and foton changed 3Hz to 0.375Hz (2048/16384) and other coefficients.

Note:
 The damping servo for SIDE has been working for every MCs. POS, PIT, YAW are the ones not working.

Next work:
 - Fix filters for every suspensions.
   There are backup files in /cvs/cds/rtcds/caltech/c1/chans/filter_archive directory, so this should help.
   But I don't think just fixing filters would fix the damping servo because the filter design of MC suspensions were changed this morning and the damping had worked until Friday.

  3741   Tue Oct 19 15:14:51 2010 JenneUpdateSUSPRM (little) update

[Jenne, Suresh]

We've aligned the guiderod and wire standoff to the PRM, each partly.  They have both been aligned to the correct distance above the scribe lines, but they have not yet been centered forward/backward along the thickness of the optic.  So, we're working on it...

  3740   Tue Oct 19 00:24:07 2010 DmassOmnistructureElectronicsMassive restocking of the 40m

I had a number of delinquent items on the sign out list from the 40m. I returned about half, and ordered replacements for most of the other half.

I put the photodiodes on the SP table, and the 560 on the electronics  bench.

  3739   Mon Oct 18 22:11:32 2010 Joonho LeeSummaryElectronicsCCD cables for input signal

Today I checked all the CCD cables which is connected input channels of the VIDEOMUX.

Among total 25 cables for output, 12 cables are type of RG59, 4 cables are type of RG58, and 9 cables are of unknown type.

The reason I am checking the cables is for replacing the cables with impedance of 50 or 52 ohm by those with impedance of 75 ohm.

After I figures out which cable has not proper impedance, I will make new cables and substitute them in order to match the impedance, which would lead to better VIDEO signal.

 

Today, I check the cables in similar way as I did the last time.

I labeled all cables connected to input channels of VIDEO MUX and disconnect all of them since last time it was hard to check every cable because of cables too entangled.

Then I checked the types of all cables and existing label which might designate where each cable is connected to.

After I finished the check, I reconnected all cables into the input channel which each of cable was connected to before I disconnected.

 

4 cables out of 25 are type of RG58 so expected to be replace with cable of type RG59.

9 cables out of 25 are of unknown type. These nine cables are all orange-colored thick cables which do not have any label about the cable characteristic on the surface.

The result of observation is as follows.

Note that type 'TBD-1' is used for the orange colored cables because all of them look like the same type of cable.

 

Channel number where its signal is coming type
1 C1:IO-VIDEO 1 MC2 TBD-1
2 FI CAMERA 59
3 PSL OUTPUT CAMERA 59
4 BS  C:1O-VIDEO 4 TBD-1
5 MC1&3 C:1O-VIDEO 5 59
6 ITMX C:1O-VIDEO 6 TBD-1
7 C1:IO-VIDEO 7 ITMY TBD-1
8 C1:IO-VIDEO 8 ETMX TBD-1
9 C1:IO-VIDEO 9 ETMY TBD-1
10 No cable is connected
(spare channel)
 
11 C1:IO-VIDEO 11 RCR 59
12 C1:IO-VIDEO RCT 59
13 MCR VIDEO 59
14 C1:IO-VIDEO 14 PMCT 59
15 VIDEO 15 PSL IOO(OR IOC) 59
16 C1:IO-VIDEO 16 IMCT TBD-1
17 PSL CAMERA 58
18 C1:IO-VIDEO 18 IMCR 59
19 C1:IO-VIDEO 19 SPS 59
20 C1:IO-VIDEO 20 BSPO TBD-1
21 C1:IO-VIDEO 21 ITMXPO TBD-1
22 C1:IO-VIDEO 22 APS1 59
23 ETMX-T 58
24 ETMY-T 58
25 POY CCD VIDEO CH25 58
26 OMC-V 59

Today I could not figure out what impedance the TBD-1 type(unknown type) has.

Next time, I will check out the orange-colored cables' impedance directly and find where the unknown output signal is sent. Any suggestion would be really appreciated.

  3738   Mon Oct 18 18:33:46 2010 KojiSummaryCOCSummary of the main mirrors & their phasemap measurement

I have made a summary web page for the 40m upgrade optics.

https://nodus.ligo.caltech.edu:30889/40m_phasemap/

I made a bunch of RoC calculations along with the phase maps we measured.
Those are also accommodated under this directory structure.

Probably.... I should have used the wiki and copy/paste the resultant HTML?

  3737   Mon Oct 18 18:00:36 2010 KojiUpdateSUSOld PRM, SRM stored, new PRM drag wiped

- Steve is working on the storage shelf for those optics.

- PRMU002 was chosen as it has the best RoC among the three.

Quote:

[Jenne, Suresh]

We've put the old PRM and SRM (which were living in a foil house on the cleanroom optical table) into Steve's nifty storage containers.  Also, we removed the SRM which was suspended, and stored it in a nifty container.  All 3 of these optics are currently sitting on one of the cleanroom optical tables.  This is fine for temporary storage, but we will need to find another place for them to live permanently.  The etched names of the 3 optics are facing out, so that you can read them without picking them up.  I forgot to note the serial numbers of the optics we've got stored, but the old optics are labeled XRM ###, whereas the new optics are labeled XRMU ###. 

Koji chose for us PRMU 002, out of the set which we recently received from ATF, to be the new PRM.  Suresh and I drag wiped both sides with Acetone and Iso, and it is currently sitting on one of the rings, in the foil house on the cleanroom optical table.

We are now ready to begin the guiderod gluing process (later tonight or tomorrow).

 

  3736   Mon Oct 18 17:16:30 2010 JenneUpdateSUSOld PRM, SRM stored, new PRM drag wiped

[Jenne, Suresh]

We've put the old PRM and SRM (which were living in a foil house on the cleanroom optical table) into Steve's nifty storage containers.  Also, we removed the SRM which was suspended, and stored it in a nifty container.  All 3 of these optics are currently sitting on one of the cleanroom optical tables.  This is fine for temporary storage, but we will need to find another place for them to live permanently.  The etched names of the 3 optics are facing out, so that you can read them without picking them up.  I forgot to note the serial numbers of the optics we've got stored, but the old optics are labeled XRM ###, whereas the new optics are labeled XRMU ###. 

Koji chose for us PRMU 002, out of the set which we recently received from ATF, to be the new PRM.  Suresh and I drag wiped both sides with Acetone and Iso, and it is currently sitting on one of the rings, in the foil house on the cleanroom optical table.

We are now ready to begin the guiderod gluing process (later tonight or tomorrow).

  3735   Mon Oct 18 15:33:00 2010 josephb, alexUpdateCDSPossibly broken timing cards

It looks as though we may have two IO chassis with bad timing cards.

Symptoms are as follows:

We can get our front end models writing data and timestamps out on the RFM network.

However, they get rejected on the receiving end because the timestamps don't match up with the receiving front end's timestamp.  Once started, the system is consistently off by the same amount. Stopping the front end module on c1ioo and restarting it, generated a new consistent offset.  Say off by 29,000 cycles in the first case and on restart we might be 11,000 cycles off.  Essentially, on start up, the IOP isn't using the 1PPS signal to determine when to start counting.

We tried swapping the spare IO chassis (intended for the LSC) in ....

# Joe will finish this in 3 days.

# Basically, in conclusion, in a word, we found that c1ioo IO chassis is wrong.

  3734   Mon Oct 18 11:22:13 2010 JenneUpdateComputersShame on people not elogging! FrameFiles backups not working.

On the one hand, SHAME ON ALL PEOPLE WHO DON'T ELOG THINGS, such as the moving of scripts directories (it was a pain to figure out that that's part of why the backup scripts are broken).  On the other hand, the moving of the scripts directories brought to light a critical problem in the backup scheme. None of the frame files have been backed up since Joe replaced fb40m with fb, on ~23 Sept (I think).

What went down:

The frame builder was replaced, and no backup script was started up on the new machine.  Sadface.  Crontab doesn't work yet on the new machine, and also the 'ssh-add' commands give an error:

controls@fb /cvs/cds/rtcds/caltech/c1/scripts/backup $ ssh-add ~/.ssh/id_rsa
No such file or directory
controls@fb /cvs/cds/rtcds/caltech/c1/scripts/backup $ ssh-add ~/.ssh/backup2PB 
No such file or directory

Thus, I know that the backup was never running on the new fb.  However, the check-er script runs on nodus, and looks at the logfile, and since there was no script running, it wasn't adding to the log file, so the last log was an "Okay, everything worked" entry.  So, the check-er script kept sending me daily emails saying that everything was okie-dokey. 

Since all of the scripts were moved (Joe said this happened on Friday, although there's no elog about it), the check-er script, and all of the rest of the backup scripts point to the wrong places (the old scripts/backup directory), so I didn't receive any emails about the backup either way (usually it at least sends a "Hey, I'm broken" email).  This clued me in that we need to check things out, and I discovered that it's all gone to hell.

Since I can't add the ssh clients to the new fb, I can't actually log in to the backup computers over in Powell-Booth to check when the last legitimately successful backup was. But I suspect it was just before the fb was replaced.

So, we need to get Crontab up and running on the new Frame Builder machine so that we can run cron jobs, and we also need to figure out this backup hullabaloo.  I think I'll email / call Dan Kozak over in downs, who was talking about upgrading our backup scheme anyway.

 

  3733   Mon Oct 18 09:01:48 2010 steveHowToGeneralHow not to

This BNC cable is crying for help. Please do not do this to me.  It should be reported  to the  abused center.Throw this cable into the garbage now.

Attachment 1: P1060926.JPG
P1060926.JPG
  3732   Mon Oct 18 08:44:36 2010 ranaUpdateCDSslow DAQ channels added

Why EXCMON? I think that we should modify the filter coefficients so that the decimation filter that makes OUT16 is a 2nd order Butterworth at 1 Hz instead of whatever bogus thing it is now. Then we can delete the EXCMON from the DAQ and add either OUTPUT or OUTMON. That way we will have a low frequency signal (OUT16) and a sort of aliased RMS (OUTMON).

  3731   Fri Oct 15 22:16:22 2010 kiwamuUpdateSUSwhat's wrong with MC2

[Yuta, Suresh, Rana and Kiwamu]

 The DC alignment problem of MC2 was fixed.

There were some loosely connected cables on the backside of a VME rack which contains the MC2 SOS driver.

We pushed those connectors to make them tightly connected.  And then the problem disappeared.

 


(voltage unbalance on coils)

 Before fixing it Yuta opened the satellite box and measured the voltage across the coils using a voltmeter.

At that time UL and LR showed 20 times smaller voltages than that of the other two when we moved the DC alignment slider from min. to max. on the medm screen.

This behavior is exactly consistent with the wired motion of a beam spot which we saw when we were aligning MC2. 

 

(diagnostic using optical lever)

 After pushing the connectors, we made an optical lever using a red laser pointer in order to check the actual motion of MC2.

We confirmed that MC2 respond correctly to the alignment slider.

Quote:

 2010:Oct.14th:

It turned out that the DC alignment of MC2 from epics doesn't helathily work.

For example, the pitch slider does drive the yaw alignment as well.

 

  3730   Fri Oct 15 21:25:23 2010 SureshUpdateIOO2W NPRO laser output power drop question

  The power meter used in the measurements of elog entries 2822, 3698 and 3709 was the Ophir PD300-3W.  This power head has several damaged patches  and a slight movement of the laser spot changes the reading considerably.  To verify I checked the power out with another power meter (the Vector S310) and found that there is no significant variation of the power output with the temperature of the laser.  And the power at 2.1A of diode current is 2W with 10% fluctuation arising from slight repositioning of the laser head.  There are regions of the Ophir PD300 which show the laser power to be about 1.9W.

  3728   Fri Oct 15 15:32:35 2010 josephbUpdateCDSslow DAQ channels added

I added all the _OUT16, _INMON, _EXCMON channels associated with the BS, ITMX, and ITMY channels in SUS_SLOW.ini in the /opt/rtcds/caltech/c1/chans/daq directory.  Similarly, I added the channels associated with MC1, MC2, and MC3 to MCS_SLOW.ini and those associated with PRM and the SRM to RMS_SLOW.ini. 

Lines pointing to these files were then added to the /opt/rtcds/caltech/c1/target/fb/master file and the frame builder restarted.  It took about 4 tries before the frame builder stayed up using ./daqd -c ./daqdrc.

I generated the .ini files with a python script.  Its at /opt/rtcds/caltech/c1/scripts/daqscripts/create_inmon_out16_daq_ini.py. It checks the C0EDCU.ini to find what slow channels already exist, then goes through a medm directory given at the command line looking for all _OUT16, _INMON, and _EXCMON channels, adding them if they aren't already accounted for, and then writing out the new file to the location of your choice.

  3727   Fri Oct 15 06:31:52 2010 kiwamuUpdateGreen LockingPSL beat note setup: part II

 I made some KAIZENs (what does kaizen mean ? ) for the PSL green setup.

I replaced the lenses for the modematching of the two green lights at the PSL table, and the beams now look pretty identical.

Also I tuned the temperature setpoint of the doubling crystal and eventually the green light increased to 14 uW at the PSL table.

Once I finish the modification of the RFPD tomorrow, I am going to search for the beat note signal.

 


( details )

 - In-vac green mirrors

   I found one of the green steering mirror, which stands at the corner of the MC table, was clipping the green light.

  So I steered another mirror, which sends the beam to the clipping mirror after the downward periscope.

 I touched also the last steering mirror in the OMC chamber to correct the alignment.

 

 - temperature of the doubling crystal

   I took a quick temperature scan in order to find an optimum point for the crystal temperature.

  The scan was performed by just turning the heater off after I heated up the crystal up to 40 deg. 

  Using the NewPort power meter I found the optimum point around 37.3 deg. So I set the temperature to that point.

 

 - mode matching lenses

 As written in this entry , Kevin and I had put some lenses to make the two green beam almost the same size at the RFPD.

 But today while I was checking these mode-profiles by using a sensor card, I found they were not so matched.

 Therefore I replaced these lenses to match them more.

The idea of this replacement is t to let them have a long Rayleigh range, such that they can efficiently and easily interfere because of the flatness of their wave fronts along a distant.

 For the green light from the chamber, I put one more lens to form a Keplerian beam shrinker (see here about the Keplerian lens configuration).

 They look pretty identical now.

 

  3726   Fri Oct 15 00:15:52 2010 KojiUpdateIOO2W NPRO laser output power versus temperature

From the plot, you observed the reduction of the output power only by 1% between 25deg to 45deg.
This does not agree with the reduction from 2.1W to 1.6W.
Is there any cause of this discrepancy?

Quote:

 

tempscan.png

 

The measurement was made by attenuating the roughly 2W laser beam by a stack of two Neutral Density filfers and then measuring the transmitted light with the PDA36A photodetector.  This was because both the power meters used in the past were found to have linear drifts in excess of 30% and fluctuations at the 10% level. 

 

  3725   Thu Oct 14 23:33:45 2010 SureshUpdateIOO2W NPRO laser output power versus temperature

Steve measured an apparent power drop in the 2W NPRO output from 2.1W to 1.6W(elog entry no 3698) at 2.1A of diode current in the laser (elog entry:  2822).  It was later noticed that the laser temperature was set to about 45 degC while the initial calibration was done at 25 deg C.  

It was felt that the recent power drop may have something to do with the increase in the operating temperature of the laser from 25 to 45 deg C.  Therefore the laser was returned to 25 deg C and the power output was remeasured and found to be 2.1W as it was at the begining(elog entry:3709)

It was also noticed that returning the laser to 25 deg. C resulted in a loss of efficiency in coupling to the PMC.  We suspected that this might be due to multimode operating conditions in the laser at particular operating temperatures.  In order to see if this is indeed the case the laser power output was observed as a function of temperature.  We do notice a characteristic saw-tooth shape which might indicate multimode operation between 39 and 43 deg C.  It is best to verify this by observing the power fluctuations in the transmitted beam of the stabilised reference cavity.

 

tempscan.png

 

The measurement was made by attenuating the roughly 2W laser beam by a stack of two Neutral Density filfers and then measuring the transmitted light with the PDA36A photodetector.  This was because both the power meters used in the past were found to have linear drifts in excess of 30% and fluctuations at the 10% level. 

 

 

Attachment 2: Scan2010.zip
  3724   Thu Oct 14 22:26:38 2010 KojiUpdateComputersNew Netgear Switch

The network cables for the Martian network were moved to the new Netgear switch from the old one which had the broken fan.
The martian machines look happy so far.

Above the new switch we have the GC network switch. The two fans of it were also broken. The fans were replaced.

They are now quiet and I am quite satisfied.

Quote:

I removed some old equipment from the rack outside the control room and stacked them next to the filing cabinets in the control room. I also mounted the new Netgear switch in the rack.

 

  3723   Thu Oct 14 20:49:50 2010 yutaUpdateComputersautomated Q-value adjuster

Background:
 Now we can use pyNDS(see elog #3722), so I wrote a python script that adjusts Q-value automatically.

What I did:
1. Wrote a python script. /cvs/cds/caltech/users/yuta/scripts/QAdjuster.py

2. Ran this script for every DOF, every MC suspension.

Result:
 Following lines and attached are example output when I ran QAdjuster.py for MC3 POS.

Currently, C1:SUS-MC3_SUSPOS_GAIN is 5

Connecting.... authenticate failed: Unspecified error
Connecting.... done
Current GPS time is 971149410

I kicked C1:SUS-MC3_SUSPOS_OFFSET

Waiting for the ringdown...... 0\

Current GPS time is 971149431

omega_0= 6.326764,
Q= 11.254252,
A= -53.491850,
delta= 5.007821,
ofset= 4349.808434

Q-value was 11.3
Set C1:SUS-MC3_SUSPOS_GAIN = 10.2311380794


 Current Q-values automatically set are as follows.

  MC1 MC2 MC3
POS 4.9 4.8 5.9
PIT 6.8 5.5 4.5
YAW 6.4 5.6 5.3
SIDE 5.8 5.2 5.0


Next work:
 - Modify the script so that it adjusts all the channels automatically. Now, it is one by one, trial by trial.
 - Modify the script so that it automatically turns the offset switch on. Now, it must be turned on beforehand.
 - Write a how-to

Attachment 1: MC3_SUSPOS.png
MC3_SUSPOS.png
  3722   Thu Oct 14 20:31:15 2010 yutaUpdateComputerspyNDS installation

EDIT by kiwamu Dec.7th ;

The Pynds package has been moved to the appropriate place:

/cvs/cds/caltech/apps/linux64/python/lib64/python2.4/site-packages

Background:
 We need the python module pyNDS to get the data directly from the server using python.
 Leo helped us install pyNDS today.

Installation:
1. Get pyNDS source tarball from here.
  https://www.lsc-group.phys.uwm.edu/daswg/download/software/source/pynds-0.3.tar.gz

2. All you need is Boost.Python and nds2-client. See README in the pynds-0.3 directory.

3. I installed pyNDS to /cvs/cds/caltech/users/yuta/pynds

4. You have to set environment variable to import the module. For example, run;
  setenv PYTHONPATH /users/yuta/pynds/lib64/python2.4/site-packages:${PYTHONPATH}

Notes:
 RPM will be available soon from Leo.

  3721   Thu Oct 14 14:52:52 2010 Leo SingerConfigurationComputersnumpy, ipython, matplotlib, python-matplotlib installed on rossa

I installed the following packages on rossa:

numpy, ipython, matplotlib, python-matplotlib

  3720   Thu Oct 14 14:09:30 2010 josephbUpdateCDSNDS 2 server status

[Joe, John]

The nds 2 server is about 50% of the way there.  You can connect to it and get channel lists, but its having issues actually serving the data.  The errors we're getting basically say it can't find the source data for the channel

John had to go get lunch at 2pm, but he said he'd log in remotely later and try to configure it properly.

  3719   Thu Oct 14 13:15:14 2010 Leo SingerConfigurationComputersgit installed on rossa

I installed git on rossa using:

$ sudo yum install git

  3718   Thu Oct 14 13:09:01 2010 Leo SingerConfigurationComputersnds2-client-devel installed on rossa

I installed nds2-client-devel on rossa using the following command:

$ sudo yum install nds2-client-devel

  3717   Thu Oct 14 12:53:29 2010 yutaUpdatePSLmesured PMC's visibility vs power relation

Background:
 I measured the PMC's visibility vs incident power relation last week to see the thermal effect, but I didn't calibrated the laser power(see elog #3672).
 So, I calibrated it on Oct 12.

Setup:

 Attachment #1
 The laser grade window PW-1025-UV-1064-45P had power reflectivity of about 0.5% in this setup.

What I did:
1. Calibrated the laser power(Attachment #2).
  To measure the laser power, I put the Ophir power meter at just in front of "PMC REFL PD".

2. For the calculation of the visibility K, I used the following formula;
 K= [1-(R1-R0)/(R2-R0)]*100
where R0, R1 and R2 are the PD outputs in voltage when laser is off, PMC locked and not locked respectively.

3. Plotted the visibility vs the incident power(Attachment #3).

Result:
 Attachment #2
  From the linear fit by least squares, the calibration turned out to be 1.12±0.07 mV/uW. The error of this value is calculated from assuming PD output error~1mV and laser power error~3uW for all measured value.
  The largest error was from the position and the angle of the power meter probe.

 Attachment #3
  I used the same data I took last week(see elog #3672), but better plot.
  I put the error bars for just several points. When the laser power is weak, the errors are large because of the cancellation error. When the laser power is high, the errors are estimated to be so small that you can't see it in the plot(~1%).
  At the several points, I couldn't lock the PMC well and  the power of the reflected light depended on the offset voltage of the PZT.
  The horizontal axis has about 6% error because of the calibration error.

Note:
 Now the condition is a bit different from this measurement(NPRO temperature changed, optics moved slightly), so the visibility may be changed.

Attachment 1: PMCsetup.png
PMCsetup.png
Attachment 2: PDcalib.png
PDcalib.png
Attachment 3: PMCreflect2.png
PMCreflect2.png
  3716   Thu Oct 14 12:45:47 2010 josephbUpdateComputersNDS 2 server installation on mafalda

[Joe, John Zweizig]

John stopped by around noon today to install the NDS 2 server.  He installed it /cvs/cds/caltech/users/jzweizig/nds2-server/.

Once John is done, I will be moving this to a more sensible install location that is not his user directory, but its there for the moment.

We had to install a couple more packages including bzip2, bzip2-devel, gcc-c++, openssl, and openssl-devel. 

We mounted the /frames directory from the fb machine to mafalda by modifying the /etc/fstab file with the line:

fb:/frames              /frames                 nfs     bg,ro           0 0

 

If we change channels recorded by the frame builder, we need to update a channel list file for the NDS 2 server.  There's an excutable located at:

/cvs/cds/caltech/users/jzweizig/nds2-server/bin/buildChannelList

This builds the file list if given a .gwf file.  These are written by the frame builder, and can be found in /frames/full/####, where #### are the first 4 gps digits of the gravity files contained in that directory.

Upon questing about when we get to GPS time 1000000000, he said there's some updates he needs to do so it rather throws away the last 5 digits, rather than keeping the first 4.

An example command run on the fb or mafalda machine is:

/cvs/cds/caltech/users/jzweizig/nds2-server/bin/buildChannelList  /frames/full/9711/C-R-971119728-16.gwf > nds2-mafalda/C-R-ChanList.txt

For a seconds trends file (located in /frames/trends/second/ instead of /frames/full)

/cvs/cds/caltech/users/jzweizig/nds2-server/bin/buildChannelList  /frames/trends/second/9711/ C-T-971106780-60.gwf > nds2-mafalda/C-T-ChanList.txt

For a minute trends file (located in /frames/trends/minute)

/cvs/cds/caltech/users/jzweizig/nds2-server/bin/buildChannelList  /frames/trends/minute/9711/C-T-971106780-60.gwf  > nds2-mafalda/C-M-ChanList.txt

In these cases, John was putting the lists in the /cvs/cds/caltech/users/jzweizig/nds2-mafalda/ directory.

 

Both the  C-raw-cache.txt file and the nds2.conf files need to be configured to point at the correct files in the nds2-mafalda directory.

 

  3715   Thu Oct 14 10:59:10 2010 josephbUpdateComputersUpdated cshrc.40m and Computer Restart Procedures

I started modifying the cshrc.40m file in /cvs/cds/caltech/ so that it starts pointing at the new directories.

# misc aliases
alias target 'cd /opt/rtcds/caltech/c1/target'
alias scripts 'cd /opt/rtcds/caltech/c1/scripts'
alias chans 'cd /opt/rtcds/caltech/c1/chans'
alias c 'cd /opt/rtcds/caltech/c1'
alias s 'cd /opt/rtcds/caltech/c1/scripts'
alias u 'cd /cvs/cds/caltech/users'

I also added "alias screens 'cd /opt/rtcds/caltech/c1/medm'" as a quick way to get to the medm directory.

Once we get multiple compiled versions (i.e. i386, x86_64, Solaris) of the new gds tools from Alex, we'll have to some more serious surgery on this file.

I removed the "New Computer Restart Procedures" and simply moved the changes into the "Computer Restart procedures", found here.  I've removed everything I don't think applies anymore (all the VME FE reboot procedures for example).

  3714   Thu Oct 14 10:29:33 2010 josephbUpdateComputersMafalda ready for NDS2, updated Rosalba, Rossa repos

At John Zweizig's request, I installed a couple of packages from the lscsoft repository, along with libtools, automake, autoconf, and several kerberos packages, including cyrus-sasl, cyrus-sasl-lib, cyrus-sasl-devel, cyrus-sasl-gssapi, and krb5-libs.  All it needs now is John to come down and install the NDS2 server.

I copied the lscsoft.repo file from /etc/yum.conf.d/ on allegra to mafalda, as well as rosalba and rossa, just for completeness.  I also added the epel repository to rossa and installed the yum-priorities package and set the priorities on rossa's repositories. 

  3713   Thu Oct 14 01:09:17 2010 yutaUpdateIOOtried to lock MC but failed

(Rana, Koji, Kiwamu, Suresh, Yuta)

 We attempted to lock the MC and finally got flashes of the MC, but no luck. 

Tomorrow we are going to check every components one by one to make sure if everything is okay or not.


Background:
 MC suspensions are well damped now.
 We need MC locked for the alignment of the in-vac optics.

Issues:

  These are the issues which we are going to fix.
1. DC alignment of the MC2 suspension doesn't seem to be working correctly. (see here)

  We should check the satellite box and the cable connection.

  The coils look like woking fine because we can kick MC2 by using each of the coils.


2. Incomplete modematching.

The spot size of the reflected light from MC1 looked like bigger. 

3. beam axis of the injection light to MC1

  3712   Thu Oct 14 00:54:12 2010 ranaHowToCDSBad PSL Quad, so I edited the SUS MEDM screens

We found today that there was some (unforgivably un-elogged) PSL-POS & PSL-ANG work today.

The wedge splitter at the end of the PSL table is so terribly dogged that it actually moves around irreversibly if you touch the post a little. Pictures have been recorded for the hall of shame. This should be replaced with a non-pedestal / fork mount.

I was so sad about the fork clamp, that I edited the SUS Master screen instead according to Joe-Yuta instructions; see attached.Untitled.png

There's clearly several broken/white links on there, but I guess that Yuta is going to find out how to fix these from Joe in the morning.

  3711   Thu Oct 14 00:53:44 2010 kiwamuUpdateSUSwhat's wrong with MC2

It turned out that the DC alignment of MC2 from epics doesn't helathily work.

For example, the pitch slider does drive the yaw alignment as well.

 

Is this somehow related to the unknown MC2 jump happened around September 10th ?? (see the trend below)

OSMEs.png

  3710   Thu Oct 14 00:04:46 2010 yutaUpdateSUSQ-value adjustment for MC dampings

Background:
 We need MC to be locked soon, so MC suspensions should be damped well(Q~5).

What I did:
1. Set the correct filters and turn all the damping servo on.

2. Kick the optics by adding some offset into the control loop.

3. Measure the Q-value of the ringdown with my eye.

4. If Q-value seems to be around 5, go to step #5. If not, change the filter gain and go to step #2.

5. Done! Do step #1-4 for all MCs.

All parameters I used for the servo are automatically saved here;
  /cvs/cds/caltech/burt/autoburt/snapshots/2010/Oct/13/20:07/c1mcs.epics

Result:
 Q-values of the damping servo for all MCs are set to around 5.
 Attached is the ringdown of MC2 for example.
 As you can see, my eye was very rough......

Next work:
 - Make a script that does steps #2-5 automatically.
    I need pyNDS module installed to get data using Python.
    I already wrote the rest of the script.
    We'll have Leo help us install pyNDS tomorrow.

Attachment 1: MC2ringdown.png
MC2ringdown.png
  3709   Wed Oct 13 21:08:40 2010 kiwamuUpdatePSLNPRO is still alive

 The NPRO at the PSL table still can generate 2W laser !  He is still alive.

  When I reduced the temperature to  25 deg, the output power increased to 2W successfully.

  As Steve wrote down in his last entry (see here), the NPRO output was at 1.6 W currently, which is supposed to be 2W.

We were suspicious about the laser crystal's temperature because the current temperature looks a bit high.

In fact the setpoint of the temperature was 45.9 deg instead of 25 deg that is the previous setpoint.

 

ELOG V3.1.3-